the perimeter of certain rectangule is16 timesthe width. the lengthis 12cm. more than the width. find the widthodf the rectangule
Answer by nerdybill(2692)

the perimeter of certain rectangule is16 timesthe width. the lengthis 12cm. more than the width. find the widthodf the rectangule
.
Let w = width of the rectangle
then
w+12 = length of the rectangle
.
perimeter = 2(width + length)
16w = 2(w + w+12)
16w = 2(2w+12)
16w = 4w+24
12w = 24
w = 2 cm (width)